# UniHR: Hierarchical Representation Learning for Unified Knowledge Graph Link Prediction

### v3

@@ Abstract
Beyond-triple fact representations including hyper-relational facts with auxiliary key-value pairs, temporal facts with additional timestamps, and nested facts implying relationships between facts, are gaining significant attention. However, constrained by complex fact representation forms, existing link prediction models for beyond-triple facts have difficulty achieving hierarchical fact modeling and generalizing the modules for one specific facts to other fact types. To overcome this limitation, we propose a Unified Hierarchical Representation learning framework (UniHR) for unified knowledge graph link prediction. It consists of a unified Hierarchical Data Representation (HiDR) module and a unified Hierarchical Structure Learning (HiSL) module as graph encoder. The HiDR module unifies hyper-relational KGs, temporal KGs, and nested factual KGs into triple-based representations. Then HiSL incorporates intra-fact and inter-fact message passing, focusing on enhancing the semantic information within individual facts and enriching the structural information between facts. Empirical results demonstrate the effectiveness of UniHR and highlight the strong potential of unified representations. Code and data are available at https://github.com/Lza12a/UniHR.
Keywords:
Knowledge graph representation learning Beyond-triple knowledge graph Link prediction.
1 Introduction
Large-scale knowledge graphs (KGs) such as Freebase [3] and Wikidata [33] have been widely applied in many areas like question answering [17] and natural language processing [2]. However, the presence of missing facts within these KGs inevitably limit their applications. Therefore, the link prediction task has been introduced to predict missing elements within factual data. Current link prediction methods mainly focus on the fact in the form of triple e.g., $(head\,entity,relation,$
$tail\,entity)$.
@@ 1 Introduction
Figure 1: Comparison between the existing link prediction methods for specific beyond-triple knowledge graphs and our unified methods.
@@
Despite the simplicity and unity of triple-based representation, it is difficult to adequately express complex facts, such as “Oppenheimer is educated at Harvard University for a bachelor degree in chemistry" shown in Figure 1. Therefore, existing researches [34, 37, 38] contribute to focusing on semantically richer facts. Figure 1 illustrates three specific types of facts: hyper-relational fact ((Oppenheimer, educated at, Harvard University), degree: bachelor, major: chemistry), temporal fact (Oppenheimer, honored with, Fermi Prize, 1963), nested fact ((Oppenheimer, born in, New York), imply, (Oppenheimer, nationality, The United States)). These forms of facts allow for expression of complex semantics and revelation of relationships between facts, extending beyond the triple-based representation. Thus in recent years, Hyper-relational KGs (HKG) [6], Temporal KGs (TKG) [40], and Nested factual KGs (NKG) [37] attract wide research interests.
Recent studies have demonstrated the effectiveness of various embedding strategies for these beyond-triple representations [36]. However, these methods are hard to achieve hierarchical fact modeling and usually designed for specific representation forms, e.g., StarE [12] customizes graph neural network to enhance inter-fact structure information between hyper-relational facts. For NKGs, BiVE [7] bridges intra-fact semantics between two levels of facts through a simple MLP. Moreover, GeomE+ [40] et al. TKG methods employ time-aware scoring functions to capture intra-fact collaborative information between factual semantics and temporal information in the timestamp. Although these methods perform well on specific types of facts, it is evident that such customized methods are difficult to generalize to other types of KGs for intra-fact structure and inter-fact semantic enhancements. Therefore, establishing a unified hierarchical representation learning method for multiple types of KGs is worth to investigate.
@@
To overcome the challenges mentioned above, we propose a Unified Hierarchical Representation learning method (UniHR), which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module as the graph encoder. HiDR module standardizes hyper-relational facts, nested factual facts, and temporal facts into the form of triples without loss of information. Furthermore, HiSL module captures local semantic information during intra-fact message passing and then utilizes inter-fact message passing to enrich the global structure information to obtain better node embeddings based on HiDR form. Finally, the updated embeddings are fed into decoders for link prediction. Empirical results demonstrate the UniHR achieves state-of-the-art or competitive performance on most metrics and highlight the strong potential of unified representations. Our contributions can be summarized as follows.
  • 1.
@@
    We emphasize the value of investigating unified KG representation method, including unified symbolic representation and unfied representation learning method for different KGs.
  • 2.
@@
  • 3.
    We conduct link prediction experiments on 9 datasets across 5 types of KGs. Compared to methods designed for one kind of KG, UniHR achieves the best or competitive results, verifying strong generalization capability.
2 Preliminaries
In this section, we introduce the definition of four types of existing knowledge graphs (KGs): triple-based KG, hyper-relational KG, nested factual KG and temporal KG, along with link prediction tasks on these types of KGs.
@@ 2 Preliminaries
A triple-based KG $\mathcal{G}_{KG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ represents facts as triples, denoted as =$\mathcal{F}\,\text{=}\{\left(h,r,t\right)|h,t\in\mathcal{V},$ $r\in\mathcal{R}\}$, where $\mathcal{V}$ is the set of entities and $\mathcal{R}$ is the set of relations. The link prediction on triple-based KGs involves answering a query $\left(h,r,?\right)$ or (?$,r,t\right)$, where the missing element ‘????’ is an entity in $\mathcal{V}$.
@@
A hyper-relational KG (HKG) $\mathcal{G}_{HKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ consists of hyper-relational facts, abbreviated as H-Facts, denoted as =$\mathcal{F}\,\text{=}\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})$ $|\,h,t,v_{i}\in\mathcal{V},r,k_{i}\in\mathcal{R}\}$. Typically, we refer to $\left(h,r,t\right)$ as the main triple in the H-Fact and $\left\{\left(k_{i}\text{:}\,v_{i}\right)\right\}_{i=1}^{m}$ as $m$ auxiliary key-value pairs. Similar to link prediction on triple-based KGs, the link prediction on HKGs aims to predict entities in the main triple or the key-value pairs. Symbolically, the aim is to predict the missing element, denoted as ‘?’ for queries $((h,r,t),(k_{1}\text{:}\,v_{1}),\ldots(k_{i}\text{:}\,?)),$ ((?$,r,t),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$ or $((h,r,?),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$.
@@
A nested factual KG (NKG) can be represented as $\mathcal{G}_{NKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\hat{\mathcal{R}},\hat{\mathcal{F}}\}$, which is composed of two levels of facts, called atomic facts and nested facts. =$\mathcal{F}\,\text{=}\left\{\left(h,r,t\right)|h,t\in\mathcal{V},r\in\mathcal{R}\right\}$ is the set of atomic facts, where $\mathcal{V}$ is a set of atomic entities and $\mathcal{R}$ is a set of atomic relations. ^=$\hat{\mathcal{F}}\,\text{=}\,\{\left(\mathcal{F}_{i},\hat{r},\mathcal{F}_{j}\right)|$ $\mathcal{F}_{i},\mathcal{F}_{j}\in\mathcal{F},\hat{r}\in\hat{\mathcal{R}}\}$ is the set of nested facts, where $\hat{\mathcal{R}}$ is the set of nested relations. The link prediction on the NKGs is performed on the atomic facts or nested facts. We refer to the link prediction on atomic facts as Base Link Prediction, and the link prediction on nested facts as Triple Prediction. For base link prediction, given a query $\left(h,r,?\right)$ or (?$,r,t\right)$, the aim is to predict the missing atomic entity ‘????’ from $\mathcal{V}$. For triple prediction, given a query (?$,\hat{r},\mathcal{F}_{j}\right)$ or $\left(\mathcal{F}_{i},\hat{r},?\right)$, the aim is to predict the missing atomic fact ‘????’ from $\mathcal{F}$.
@@
A temporal KG (TKG) $\mathcal{G}_{TKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\mathcal{T}\}$ is composed of quadruple-based facts, which can be represented as =$\mathcal{F}\,\text{=}\,\{(h,r,t,$
$[\tau_{b},\tau_{e}])|h,t\in\mathcal{V},r\in\mathcal{R},\tau_{b},\tau_{e}\in\mathcal{T}\}$, where $\tau_{b}$ is the begin time, $\tau_{e}$ is the end time, $\mathcal{V}$ is the set of entities, $\mathcal{R}$ is the set of relations and $\mathcal{T}$ is the set of timestamps. The link prediction on TKGs aims to predict missing entities ‘????’ in $\mathcal{V}$ for two types of queries (?$,r,t,[\tau_{b},\tau_{e}]\right)$ or $\left(h,r,?,[\tau_{b},\tau_{e}]\right)$.
3 Related Works
@@
Earlier HKG representation learning methods e.g., m-TransH [35], RAE [44] have generalized the triple-based approach to HKG and loosely represent the combinations of key-value pairs. Galkin et al. first customize StarE [12] based on CompGCN [31] for H-Facts to capture global structure information between H-Facts in the message passing stage, and achieves impressive results, demonstrating that the structure information of the graph in HKGs is also important. GRAN [15] introduces edge-aware bias into the vanilla transformer attention [32], while HyNT [6] designs a qualifier encoder for HKG. They both focus on intra-fact semantic dependencies but ignore the global structure information. In order to fully capture inter-fact and intra-fact information, HAHE [23] first adopts dual graph attention and edge-aware bias in the transformer attention layer for hierarchical modeling and achieves significant performance improvement. We consider the hierarchical attention mechanism of HAHE a good beginning, but it is customized for HKG and is difficult to generalize to other facts.
@@
Chung et al. [7] first introduced nested facts to imply relationships between facts. They also propose BiVE which bridges semantics between atomic facts and fact nodes in the encoding phase via a simple MLP and scores both atomic facts and nested facts using the quaternion-based KGE scoring functions like QuatE [45] or BiQUE [16]. Based on BiVE, NestE [37] represents the fact nodes as a $1\times 3$ embedding matrix and the nested relations as a $3\times 3$ matrix to avoid information loss, embedding them into hyperplanes with different dimensions. These methods only capture the intra-fact semantic information between atomic facts and nested facts. Meanwhile, due to the complexity of this representation, common triple-based GNNs have difficulty in message passing between atomic fact and nested fact for capturing inter-fact global information.
@@
Recent studies in temporal knowledge graph representation learning have focused on designing special time-aware scoring functions for inter-fact semantic modeling. Models such as TTransE [21], HyTE [8], TeRo [39], and TGeomE+ [40] incorporate temporal-aware module into the KGE score function in various ways. However, existing models seldom regard timestamps as node and directly utilize inter-fact message passing mechanism to perceive time information for enhancing entity and relation embeddings. Therefore, we believe it is a direction worth exploring.
4 Methodology
@@
In this section, we introduce our method, a Unified Hierarchical Representation learning framework (UniHR), which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module. Our workflow can be divided into the following three steps: 1) Given a KG $\mathcal{G}$ of any type, we represent it into $\mathcal{G}^{HiDR}$ under the HiDR form. 2) The $\mathcal{G}^{HiDR}$ will be encoded by HiSL module to enhance the semantic information within individual facts and structural information between facts on the whole graph. 3) In the phase of decoding, the updated embeddings of nodes and edges are fed into transformer decoders to obtain the plausibility score of facts.
Figure 2: Diverse facts are translated into HiDR form.
4.1 Hierarchical Data Representation
To overcome the differences in the representation of multiple types KGs, we introduce a Hierarchical Data Representation module, abbreviated as HiDR. Different from labelled RDF representation [1], we constrain “triple" to be considered as the basic units of HiDR form, then HiDR could continuous benefit from the model developments of triple-based KGs, which is the most active area about link prediction over KGs.
@@ 4.1 Hierarchical Data Representation
As shown in Fig. 2, in order to ensure comprehensive representation of facts, we introduce three hierarchical types of nodes and three connected relations in HiDR. Firstly, we denote original entities within three types of KGs as atomic nodes and complement fact nodes for HKGs and TKGs lacking a designated fact node. To facilitate the interaction between fact nodes and relations explicitly, we incorporate relation nodes into the graph, represented as $e_{r}$ for each $r$. These relation nodes are derived from transforming the relation edges in the original KG. It is important to facilitate direct access of fact nodes to the relevant atomic nodes during message passing process. To achieve this, we introduce three connected relations: has relation, has head entity and has tail entity, which establish directly connections between atomic nodes and fact nodes. Ultimately, we denote the (main) triple $\left(h,r,t\right)$ in original fact as three connected facts: $\left(f,has\,relation,e_{r}\right),\left(f,has\,head\,entity,h\right),$ $(f,has\,tail\,entity,t)$, and an atomic fact $(h,r,t)$, where $f$ is fact node. Formally, the definition of HiDR form is as follows:
Definition 1
@@
Hierarchical Data Representation (HiDR): A KG represented as the HiDR form is denoted as $\mathcal{G}^{HiDR}\,\text{=}\,\{\mathcal{V}^{HiDR},\mathcal{R}^{HiDR},\mathcal{F}^{HiDR}\},$ where $\mathcal{V}^{HiDR}\,\text{=}\,$
$\mathcal{V}_{a}\cup\mathcal{V}_{r}\cup\mathcal{V}_{f}$ is a joint set of atomic node set ($\mathcal{V}_{a}$), relation node set ($\mathcal{V}_{r}$), fact node set ($\mathcal{V}_{f}$). $\mathcal{R}^{HiDR}\,\text{=}\,\mathcal{R}_{a}\cup\mathcal{R}_{n}\cup\mathcal{R}_{c}$ is a joint set of atomic relation set ($\mathcal{R}_{a}$), nested relation set ($\mathcal{R}_{n}$), connected relation set $\mathcal{R}_{c}\,\text{=}\{has\,relation,has\,head\,entity,has\,$
@@
$tail\,entity\}$. The fact set $\mathcal{F}^{HiDR}\,\text{=}\,\mathcal{F}_{a}\cup\mathcal{F}_{c}\cup\mathcal{F}_{n}$ is jointly composed of three types of triple-based facts: atomic facts ($\mathcal{F}_{a}$), connected facts ($\mathcal{F}_{c}$) and nested facts ($\mathcal{F}_{n}$), where $\mathcal{F}_{a}\text{=}\,\{(v_{1},r,$ $v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{a},r\in\mathcal{R}_{a}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1}\in\mathcal{V}_{f},r\in\mathcal{R}_{c},v_{2}\in\mathcal{V}_{a}\}$, $\mathcal{F}_{n}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{f},r\in\mathcal{R}_{n}\}$.
Next, we introduce how to transform different types of KGs into HiDR form.
For hyper-relational knowledge graphs, we regard key-value pairs as complementary information for facts. Thus, we translate the H-Facts $\mathcal{F}_{HKG}\,\text{=}$
$\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m}\}$ into the HiDR form that $\mathcal{G}^{HiDR}_{HKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{HKG}\}$ following the definition, where $\mathcal{F}_{c}\,\text{=}\,\{(f,has\,relation,e_{r}),(f,has$ $\,head\,entity,h),(f,has$
$tail\,entity,t),(f,k_{1},v_{1}),\ldots,(f,k_{m},v_{m})\}$, $\mathcal{F}_{a}\,\text{=}\,\{(h,r,t)\,|\,((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})\in\mathcal{F}_{HKG})\}$ and $\mathcal{F}_{n}\text{=}\,\varnothing$.
For nested factual knowledge graphs, HiDR can naturally represent hierarchical facts, so we translate the atomic facts $\mathcal{F}_{NKG}\,\text{=}\,\{(h_{i},r_{i},t_{i})\}$ and the nested facts ^$\hat{\mathcal{F}}_{NKG}\,\text{=}\,\{((h_{1},r_{1},t_{1}),R,(h_{2},$ $r_{2},t_{2}))|(h_{i},r_{i},t_{i})\in\mathcal{F}_{NKG}\}$ into the form of HiDR that $\mathcal{G}^{HiDR}_{NKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{NKG}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\,\{(h_{i},r_{i},t_{i})\,|\,(h_{i},r_{i},t_{i})\in$ $\mathcal{F}_{NKG}\}$, $\mathcal{F}_{c}\,\text{=}\{(f_{i},has\,head\,entity,h_{i}),(f_{i},has\,tail$
$entity,t_{i}),(f_{i},has\,relation,e_{r_{i}})|f_{i}\,\text{=}\,(h_{i},r_{i},$ $t_{i})\in\mathcal{F}_{NKG}\}$ and $\mathcal{F}_{n}\text{=}\{(f_{1},R,f_{2})|f_{i}\in\mathcal{F}_{NKG}\}$.
For temporal knowledge graphs, we regard the TKG as a special HKG, and convert timestamps to auxiliary key-value pairs in HKGs by adding two special atomic relations: begin and end, regarding timestamps as special numerical atomic nodes. Thus, we firstly translate the temporal facts in TKGs $\mathcal{F}_{TKG}\text{=}\,\{(h,r,t,[\tau_{b},\tau_{e}])\}$ into H-Facts form $\mathcal{F}_{TKG}^{HKG}\text{=}\,\{(h,r,t,begin\text{:}\tau_{b},end\text{:}\tau_{e})\}$. Then according to the previous transformation in HKG, it can be translated into the HiDR form that $\mathcal{G}_{TKG}^{HiDR}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}_{TKG}^{HiDR}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\{(h,r,t)\,|\,(h,r,t,begin\text{:}\,\tau_{b},end\text{:}$ $\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(f,has\,relation,e_{r}),(f,$ $has\,head\,entity,h),\,(f,has\,tail\,entity,t),\,(f,$ $begin,\tau_{b}),(f,end,\tau_{e})\,|\,f\,\text{=}\,(h,r,t,$
$begin\text{:}\,\tau_{b},end\text{:}\,\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$ and $\mathcal{F}_{n}\,\text{=}\,\varnothing$.
In summary, we could convert all above KGs into the HiDR form, and preserve the semantics in the original KGs without loss of information.

### v4

@@ Abstract
Real-world knowledge graphs (KGs) contain not only standard triple-based facts, but also more complex, heterogeneous types of facts, such as hyper-relational facts with auxiliary key-value pairs, temporal facts with additional timestamps, and nested facts that imply relationships between facts. These richer forms of representation have attracted significant attention due to their enhanced expressiveness and capacity to model complex semantics in real-world scenarios. However, most existing studies suffer from two main limitations: (1) they typically focus on modeling only specific types of facts, thus making it difficult to generalize to real-world scenarios with multiple fact types; and (2) they struggle to achieve generalizable hierarchical (inter-fact and intra-fact) modeling due to the complexity of these representations. To overcome these limitations, we propose UniHR, a Unified Hierarchical Representation learning framework, which consists of a learning-optimized Hierarchical Data Representation (HiDR) module and a unified Hierarchical Structure Learning (HiSL) module. The HiDR module unifies hyper-relational KGs, temporal KGs, and nested factual KGs into triple-based representations. Then HiSL incorporates intra-fact and inter-fact message passing, focusing on enhancing both semantic information within individual facts and enriching the structural information between facts. To go beyond the unified method itself, we further explore the potential of unified representation in complex real-world scenarios, including joint modeling of multi-task, compositional and hybrid facts. Extensive experiments on 9 datasets across 5 types of KGs demonstrate the effectiveness of UniHR and highlight the strong potential of unified representations.
Introduction
Real-world large-scale knowledge graphs (KGs) such as Wikidata (Vrandečić and Krötzsch 2014) and DBpedia (Lehmann et al. 2015) have been widely applied in many areas including question answering (Kaiser, Saha Roy, and Weikum 2021) and natural language processing (Annervaz, Chowdhury, and Dukkipati 2018; Liu et al. 2025a). To faithfully represent complex real-world knowledge, these KGs usually incorporate not only standard triple-based facts, but also more complex and heterogeneous types of facts such as hyper-relational, temporal and nested facts.
@@ Introduction
Figure 1: Comparison between the existing link prediction methods for specific beyond-triple KGs and our unified methods for more realistic KGs with multiple fact types.
@@
Despite the simplicity of triple-based representation (i.e., $(head,relation,tail)$), such forms struggle to capture the complexity of real-world facts, e.g., “Oppenheimer is educated at Harvard University for a bachelor degree in chemistry”. Consequently, recent studies (Xiong et al. 2023b) have focused on semantically richer beyond-triple facts, including: hyper-relational fact ((Oppenheimer, educated at, Harvard University), degree: bachelor, major: chemistry), temporal fact (Oppenheimer, honored with, Fermi Prize, 1963), nested fact ((Oppenheimer, born in, New York), imply, (Oppenheimer, nationality, The United States)). These fact types allow for expression of complex semantics and revelation of relationships between facts. Thus in recent years, Hyper-relational KGs (HKG) (Luo et al. 2023), Temporal KGs (TKG) (Zhang et al. 2025a), and Nested factual KGs (NKG) (Li et al. 2025) attract wide research interests.
As shown in Figure 1, we find that existing research mainly suffers from two major limitations: (1) They fail to reflect real-world scenarios with multiple heterogeneous fact types (Xiong et al. 2023b), instead artificially dividing and only modeling a single KG type; (2) Earlier triple-based studies (Chen et al. 2021) have demonstrated the effectiveness of hierarchical fact semantic modeling (inter-fact and intra-fact). But due to the complexity of beyond-triple representations, they struggle to achieve comprehensive hierarchical semantic modeling, even generalizing to other fact types. Specifically, for HKGs, StarE (Galkin et al. 2020) customizes GNN to enhance inter-fact interactions, while GRAN (Wang et al. 2021) designs attention variants with edge-bias to capture intra-fact heterogeneity. For NKGs, NestE (Xiong et al. 2024) et al. NKG methods connect bi-level facts and only score intra-fact semantics. For TKGs, either by explicitly incorporating temporal information into the score function like GeomE+ (Xu et al. 2023), or by unfolding entity neighborhood subgraphs along temporal chain and capturing inter-fact semantics to model temporal information like ECEformer (Fang et al. 2024). Although advanced methods like HAHE (Luo et al. 2023) begin to capture hierarchical semantics for HKGs, their heterogeneity in representation limits their scalability to other fact types. Therefore, establishing a unified hierarchical representation learning method for real-world KG with multiple fact types is worth investigating.
@@
To fill this research gap, we propose UniHR, a Unified Hierarchical Representation learning framework, which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module. HiDR module standardizes hyper-relational facts, nested factual facts, and temporal facts into the form of triples without loss of information. Furthermore, HiSL module captures local semantic information during intra-fact message passing and then utilizes inter-fact message passing to enrich the global structure information to obtain better node embeddings based on HiDR form. Finally, the updated embeddings are fed into decoders for link prediction. Apart from the unification of method itself, UniHR’s unified representation enables flexible extensions. Unlike previous KG-specific models, UniHR accommodates more complex scenarios, such as compositional knowledge graphs, multi-task learning, and joint training on hybrid facts, thereby paving the way for pre-trained models across diverse KG types. Empirical results demonstrate the UniHR achieves competitive performance on most metrics without excessive increase in computational cost and highlight the strong potential of unified representations: Our contributions can be summarized as follows.
@@
  • We emphasize the value of investigating unified KG representation learning method, including unified symbolic representation and unfied learning for different KGs.
@@
  • We conduct link prediction experiments on 9 datasets across 5 types of KGs. Compared to KG-specific methods, UniHR achieves the best or competitive results, verifying strong generalization capability.
Preliminaries
@@
A triple-based KG $\mathcal{G}_{KG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ represents facts as triples, denoted as $\mathcal{F}\,\text{=}\{\left(h,r,t\right)|h,t\in\mathcal{V},$ $r\in\mathcal{R}\}$, where $\mathcal{V}$ is the set of entities and $\mathcal{R}$ is the set of relations. The link prediction on triple-based KGs involves answering a query $\left(h,r,?\right)$ or $\left(?,r,t\right)$, where the missing element ‘??’ is an entity in $\mathcal{V}$.
@@
A hyper-relational KG (HKG) $\mathcal{G}_{\text{HKG}}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ consists of hyper-relational facts (H-Facts) $\mathcal{F}$, denoted as $\mathcal{F}\,\text{=}\{((h,r,t),$ $\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})|\,h,t,v_{i}\in\mathcal{V},r,k_{i}\in\mathcal{R}\}$. Typically, we refer to $\left(h,r,t\right)$ as the main triple and $\left\{\left(k_{i}\text{:}\,v_{i}\right)\right\}_{i=1}^{m}$ as $m$ auxiliary key-value pairs. Link prediction on HKGs aims to predict entities in the main triple or the key-value pairs. Symbolically, the aim is to predict the missing element, denoted as ‘?’ for queries $((h,r,t),(k_{1}\text{:}\,v_{1}),\ldots(k_{i}\text{:}\,?)),$ $((?,r,t),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$ or $((h,r,?),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$.
@@
[…]
@@
A temporal KG (TKG) $\mathcal{G}_{\text{TKG}}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\mathcal{T}\}$ is composed of quadruple-based facts, represented as $\mathcal{F}\,\text{=}\,\{(h,r,t,[\tau_{b},\tau_{e}])|h,t\in\mathcal{V},$ $r\in\mathcal{R},\tau_{b},\tau_{e}\in\mathcal{T}\}$, where $\tau_{b}$ is the begin time, $\tau_{e}$ is the end time, $\mathcal{V}$ is the set of entities, $\mathcal{R}$ is the set of relations and $\mathcal{T}$ is the set of timestamps. The link prediction on TKGs aims to predict missing entities ‘??’ in $\mathcal{V}$ for two types of queries $\left(?,r,t,[\tau_{b},\tau_{e}]\right)$ or $\left(h,r,?,[\tau_{b},\tau_{e}]\right)$.
Related Works
This section discusses the limitations of existing methods in hierarchical modeling. See Appendix C for more discussion.
@@
Early HKG methods typically focus on modeling either local or global information. Galkin et al. (2020) customize StarE based on CompGCN (Vashishth et al. 2019) for H-Facts to capture global structure information among H-Facts, demonstrating the importance of structure information in HKGs. GRAN (Wang et al. 2021) with edge-aware bias in attention (Vaswani et al. 2017) layer, HyNT (Chung, Lee, and Whang 2023) with qualifier-aware encoder, and ShrinkE with relation-specific box (Xiong et al. 2023a) all focus on modeling intra-fact semantic information. Recent advanced methods aim to comprehensively capture both inter-fact and intra-fact information. For example, HAHE (Luo et al. 2023) employs dual-graph attention and edge-aware bias in the transformer attention layer for hierarchical modeling, achieving significant performance improvements. Similarly, HyperSAT (Wang, Chen, and Zhang 2025) accomplishes this through a combination of graph sampling and a key-value joint attention mechanism. We consider hierarchical modeling of KGs to be a promising direction, existing approaches are customized for HKG form and difficult to generalize to other types of facts.
Figure 2: Diverse beyond-triple facts are translated into the hierarchical data representation (HiDR) form.
@@
Chung et al. (Chung and Whang 2023) are the first to introduce nested facts to model relationships between facts, and also propose BiVE which bridges semantics between atomic facts and fact nodes via a simple MLP and scores both atomic facts and nested facts using quaternion-based KGE scoring functions like QuatE (Zhang et al. 2019) or BiQUE (Guo and Kok 2021). Based on BiVE, NestE (Xiong et al. 2024) represents fact nodes using a $1\times 3$ embedding matrix and the nested relations as a $3\times 3$ matrix to avoid information loss. GRADATE (Li et al. 2025) enhances entity and fact representation learning by mining latent intra-fact semantics. However, due to the complexity of NKG representations, existing methods have so far primarily focused on capturing intra-fact semantic information.
@@
Recent studies on TKG representation learning have mainly focused on designing elegant time-aware modules to enhance representation capability. Advanced models like TGeomE+ (Xu et al. 2023) improve the modeling of local semantics in TKGs through 4th-order tensor factorization and linear temporal regularization. Similarly, HGE (Pan et al. 2024) and 5EL (Zhang et al. 2025a) enhance the expressiveness of temporal spaces by introducing geometric spaces. In contrast, ECEformer (Fang et al. 2024) captures only inter-fact semantics by unfolding entity neighborhood subgraphs along the temporal chain to model temporal information. However, TKG representation learning methods that can simultaneously capture both intra-fact temporal semantics and global structural information remain largely unexplored. UniHR achieves this by regarding timestamps as nodes and directly employing hierarchical message passing.
Methodology
@@
In this section, we introduce UniHR, a Unified Hierarchical Representation learning framework, which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module. Our workflow includes the following three steps: (1) Given a KG $\mathcal{G}$ of any type, we represent it into HiDR form $\mathcal{G}^{\text{HiDR}}$. (2) The $\mathcal{G}^{\text{HiDR}}$ will be encoded by HiSL module to enhance the semantic information within individual facts and structural information between facts on the whole graph. (3) In the phase of decoding, the updated node and edge embeddings are serialized and fed into the transformer to optimize the model.
Hierarchical Data Representation
To overcome the limitations of beyond-triple representations in hierarchical modeling, we introduce a Hierarchical Data Representation module (HiDR), which is optimized for representation learning.
Unlike existing triple-based systems (Ali et al. 2022) including RDF (triple) reification, RDF-star and labeled RDF representation techniques, we constrain “triple” to serve as the basic units of HiDR form, and split “nodes” and “relations” into three types respectively, making it more suitable for graph representation learning. Meanwhile, “triple” form makes HiDR could continuously benefit from the model developments of triple-based KGs, which is the most active area of research for link prediction over KGs.
@@
As shown in Figure 2, we denote the entities in original KGs as atomic nodes and abstract fact nodes for HKGs and TKGs lacking a designated fact node. To facilitate the interaction between fact nodes and relations explicitly, we incorporate relation nodes into the graph, represented as $e_{r}$ for each relation $r$. To facilitate direct access of fact nodes to the relevant atomic nodes during message passing, we also introduce three connected relations: has relation, has head entity and has tail entity, which establish directly connections between atomic nodes and fact nodes. Ultimately, we denote the (main) triple $\left(h,r,t\right)$ in original fact as three connected facts: $\left(f,has\,relation,e_{r}\right),\left(f,has\,head\,entity,h\right),(f,has\,tail\,$
$entity,t)$, and an atomic fact $(h,r,t)$, where $f$ is fact node. Formally, the definition of HiDR form is as follows:
Definition 1.
@@
Hierarchical Data Representation (HiDR): A KG represented as the HiDR form is denoted as $\mathcal{G}^{\text{HiDR}}\,\text{=}\,$
@@
$\{\mathcal{V}^{\text{HiDR}},\mathcal{R}^{\text{HiDR}},\mathcal{F}^{\text{HiDR}}\},$ where $\mathcal{V}^{\text{HiDR}}\,\text{=}\,\mathcal{V}_{a}\cup\mathcal{V}_{r}\cup\mathcal{V}_{f}$ is a joint set of atomic node set ($\mathcal{V}_{a}$), relation node set ($\mathcal{V}_{r}$), fact node set ($\mathcal{V}_{f}$). $\mathcal{R}^{\text{HiDR}}\,\text{=}\,\mathcal{R}_{a}\cup\mathcal{R}_{n}\cup\mathcal{R}_{c}$ is a joint set of atomic relation set ($\mathcal{R}_{a}$), nested relation set ($\mathcal{R}_{n}$), connected relation set $\mathcal{R}_{c}\,\text{=}\{has\,relation,has\,head\,entity,has\,tail\,entity\}$. The fact set $\mathcal{F}^{\text{HiDR}}\,\text{=}\,\mathcal{F}_{a}\cup\mathcal{F}_{c}\cup\mathcal{F}_{n}$ is jointly composed of three types of triple-based facts: atomic facts ($\mathcal{F}_{a}$), connected facts ($\mathcal{F}_{c}$) and nested facts ($\mathcal{F}_{n}$), where $\mathcal{F}_{a}\text{=}\,\{(v_{1},r,$
$v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{a},r\in\mathcal{R}_{a}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1}\in\mathcal{V}_{f},r\in\mathcal{R}_{c},v_{2}\in\mathcal{V}_{a}\}$, $\mathcal{F}_{n}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{f},r\in\mathcal{R}_{n}\}$.